In MMW class, a student’s quiz scores are 89, 92, 82, and 77. What score does the student need on the fifth quiz to have a mean quiz score of 85? All the quizzes have equal weights. Round your answer to the nearest whole number.


1
Expert's answer
2021-11-12T14:58:54-0500
"\mu=\dfrac{89+92+82+77+x}{5}=85"

"340+x=425"

"x=85"

A student’s fifth quiz scores is 85.